### v1

@@ 7 Summary, and Conclusion
In this study, we have built upon the methodology established by Dainotti et al. (2024b) and Dainotti et al. (2024c). We use a 20% larger training set and more robust formula estimation to predict the redshifts of 276 GRBs. In addition, we perform an extensive MC simulation to evaluate how our ML model will perform in the upcoming years. For the benefit of the community, we have a publicly available and free-to-use web app designed to provide redshift estimates for any GRB with plateau emission, along with additional functionalities.
@@
  • The ML results obtained in this analysis, presented in Sec. 5.1, have similar accuracy with those presented in Dainotti et al. (2024c) but slightly lower. Specifically, we have a 10% decrease in $r$, from 0.719 to 0.644, and a 9.8% increase in RMSE, from 0.91 to 1. These changes in the metrics can be attributed to two reasons. First, the new sample of GRBs added to the training set has a higher percentage of GRBs at larger redshift, as compared to the training set in Dainotti et al. (2024c) (see Sec. 3.1). This can lead to the ML model having to adjust to the new data with a larger intrinsic scatter, thus leading to a decrease in the performance metrics. The second reason is the change in the methodology for formula estimation. In Dainotti2024GRBRedshift, the formulas estimated for GAM and GLM were determined using a single training-test split instance. Here, we have expanded the methodology and generalized the formulas. Generalizing the formula here means it can better capture the underlying relations among GRB properties for a larger population of GRBs rather than just a small, limited sample of the training set. However, this generalization leads to poorer performance on a singular training-test split as the variability of the underlying relation is higher in a smaller sample. However, the formulas generalize well to additional synthetic GRB samples with a more uniform redshift distribution, which is evident from the improved balanced sampling result across all the metrics (lower right and left panels of Fig. 13). This promising result indicates that our ML model has effectively generalized to the entire GRB sample and the MC simulations further emphasize this point.
@@
  • In Sec 5.5, we presented calculations on how many GRB’s redshift estimates we will be able to obtain in the future using SVOM data, and possibly THESEUS. The numbers mentioned there are based on current estimates from publications of SVOM and Swift’s observational statistics. These numbers will need to be updated as more GRBs by SVOM are detected. Regardless, the numbers paint a promising picture of the future applicability of such ML techniques. In fact, this work has already enabled us to increase the Swift GRB sample with redshift by $\sim$110%.

### v2

@@ 6 Discussion
Another point we would like to highlight here is that even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the data required to reliably train more complex ML models, such as neural networks. Indeed, without including the plateau, Swift XRT observed 1420 GRBs up to December 2024, which is too small to train complex ML models properly. In some of our previous works, using Active Galactic Nuclei data (Dainotti et al., 2021a; Narendra et al., 2022; Gibson et al., 2022), we have shown that SuperLearner picks different, more complex models (e.g., XGBoost, EARTH) for the larger sample size. Similarly, we tested multiple ML models to find the best set of models that can give accurate redshift predictions with the current data set. The three chosen models performed the best during our analysis. Given the size of our data set, parametric and semi-parametric functions like GLM and GAM perform better. This fact is further highlighted in the Appendix in Fig. 20, where the weight assigned by SuperLearner to the model EARTH is lower than the weights assigned to the GAM, GLM, and Random forest models. This is in contrast with the results presented in Narendra et al. (2022) where Earth was one of the best models. Furthermore, to test whether the performance of these models is maintained as the dataset increases, we have also performed a Monte Carlo simulation (see Sect. 5.4). From those simulations we can say that our models will be reliable for up to 3 years, after which we will require to retrain them with different models and training data. Thus, based on our analysis, these chosen models are best suited for the data at hand.
@@ 7 Summary, and Conclusion
In this study, we have built upon the methodology established by Dainotti et al. (2024b) and Dainotti et al. (2024c). We use a 20% larger training set and more robust formula estimation to predict the redshifts of 276 GRBs. In addition, we perform an extensive MC simulation to evaluate how our ML model will perform in the upcoming years. For the benefit of the community, we have a publicly available and free-to-use web app designed to provide pseudo-redshifts for any GRB with plateau emission, along with additional functionalities.
@@
  • The ML results obtained in this analysis, presented in Sec. 5.1, have similar accuracy with those presented in Dainotti et al. (2024c) but slightly lower. Specifically, we have a 10% decrease in $r$, from 0.719 to 0.644, and a 9.8% increase in RMSE, from 0.91 to 1. These changes in the metrics can be attributed to two reasons. First, the new sample of GRBs added to the training set has a higher percentage of GRBs at larger redshift, as compared to the training set in Dainotti et al. (2024c) (see Sec. LABEL:comparesets). This can lead to the ML model having to adjust to the new data with a larger intrinsic scatter, thus leading to a decrease in the performance metrics. The second reason is the change in the methodology for formula estimation. In Dainotti2024GRBRedshift, the formulas estimated for GAM and GLM were determined using a single training-test split instance. Here, we have expanded the methodology and generalized the formulas. Generalizing the formula here means it can better capture the underlying relations among GRB properties for a larger population of GRBs rather than just a small, limited sample of the training set. However, this generalization leads to poorer performance on a singular training-test split as the variability of the underlying relation is higher in a smaller sample. However, the formulas generalize well to additional synthetic GRB samples with a more uniform redshift distribution, which is evident from the improved balanced sampling result across all the metrics (lower right and left panels of Fig. 13). This promising result indicates that our ML model has effectively generalized to the entire GRB sample and the MC simulations further emphasize this point.
@@
  • In Sec 5.5, we presented calculations on how many GRB’s pseudo-redshifts we will be able to obtain in the future using SVOM data, and possibly THESEUS. The numbers mentioned there are based on current estimates from publications of SVOM and Swift’s observational statistics. These numbers will need to be updated as more GRBs by SVOM are detected. Regardless, the numbers paint a promising picture of the future applicability of such ML techniques. In fact, this work has already enabled us to increase the Swift GRB sample with redshift by $\sim$110%.
@@ 8 Acknowledgments
9 Appendix
Here, we are presenting the histogram distributions of the training set from Dainotti et al. (2024c) and the training set used in this work. The Fig. 19 shows the previous sample with solid lines and the new sample with dashed lines. Each feature has the KS-test and AD-test statistics mentioned at the top of the histogram. As mentioned in Sec 3, $T_{90}$, and $\gamma$ obtain p-values $<$ 0.05 in both AD-test and KS-test, while $\log(F_{a})$ obtains p-value $<$ 0.05 only in AD-test. The case can be explained by looking at the distribution, where we see a heavy-tailed distribution for the newer data set (dashed lines). Since the AD-test is more sensitive to heavy-tailed distributions, we obtain a p-value of $<$ 0.05.
Figure 19: The histogram distribution comparing the current training set with the previous training set from Dainotti et al. (2024c). At the top of each distribution, the KS-test and AD-test comparison results are shown.
The weights assigned by SuperLearner to its constituent models vary depending on the data set it is being used. Fig.20 shows the weight assigned to EARTH is lower than those of Random forest, GAM, and GLM models. However, in Narendra et al. (2022), EARTH was the highest-weighted model. Thus, we can say that our final ensemble model is well-suited for the size of our data sample.
Figure 20: Barplot showing the weights assigned by SuperLearner to an ensemble of GAM, GLM, Random forest and EARTH models.